zarkam21:

2 ft 8 in + 8 ft 9 in what is the answer simplified

bm717:

First, let's add the inches. What do you get?

zarkam21:

10 ft 17 in

zarkam21:

actually this is the little sister of the owner of this account. I am in 5th grade. She told me if I need help to go ahea and post my question, someone will help me

bm717:

I see. 17 inches is really 1 feet and 5 inches, correct?

zarkam21:

yes

bm717:

So the answer is 11feet and 5 inches, because you add the 1 to the original 10.

zarkam21:

Okay

Bob:

8 + 9 in = 17 12 inches in a foot 17 - 12 = 5 2 + 8 + 1 ft = 11 11 ft and 5 in

MoonShadow:

I think the answer is 11 ft and 5in.